# On the Alexander polynomial of special alternating links

### v2

@@ Abstract.
The Alexander polynomial (1928) is the first polynomial invariant of links devised to help distinguish links up to isotopy. In recent work of the authors, Fox’s conjecture (1962) – stating that the absolute values of the coefficients of the Alexander polynomial for any alternating link are unimodal – was settled for special alternating links. The present paper is a study of the special combinatorial and discrete geometric properties that Alexander polynomials of special alternating links possess along with a generalization to all Eulerian graphs, introduced by Murasugi and Stoimenow (2003). We prove that the Murasugi and Stoimenow generalized Alexander polynomials can be expressed in terms of volumes of root polytopes of unimodular matrices, building on the beautiful works of Li and Postnikov (2013) and Tóthmérész (2022). We conjecture a generalization of Fox’s conjecture to the Eulerian graph setting. We also bijectively relate two longstanding combinatorial models for the Alexander polynomials of special alternating links: Crowell’s state model (1959) and Kauffman’s state model (1982, 2006).
Karola Mészáros received support from CAREER NSF Grant DMS-1847284 as did Elena S. Hafner and Alexander Vidinas.
@@ 1. Introduction
A tantalizing conjecture of Fox from 1962 [4] regarding the coefficients of the Alexander polynomials of alternating links remains stubbornly open to this day. It states that the absolute values of the coefficients of the Alexander polynomial of any alternating link are unimodal. For alternating links $L$, [3, 13, 14] show that $\Delta_{L}(-t)\in\mathbb{Z}_{\geq 0}[t]$, so Fox’s conjecture can be restated in the following way:
@@
[4] Let $L$ be an alternating link. Then the coefficients of $\Delta_{L}(-t)$ form a unimodal sequence.
Stoimenow [20] strengthened Fox’s conjecture to log-concavity without internal zeros. Fox’s conjecture remains open in general although some special cases have been settled by Hartley [6] for two-bridge knots, Murasugi [15] for a family of alternating algebraic links, and Ozsváth and Szabó [17] for the case of genus $2$ alternating knots. Jong [7] also confirmed that Fox’s conjecture holds for genus $2$ alternating knots. The authors of the current paper recently resolved Fox’s conjecture for special alternating links [5].
@@ 1.1. Notation and graph theory definitions
For any graph $G$, let $V(G)$ denote the vertex set of $G$ and $E(G)$ the edge set of $G$. For a digraph $D$, we will denote the initial vertex of any $e\in E(D)$ by $\operatorname{init}(e)$ and the final vertex of $e$ by $\operatorname{fin}(e)$. The number of edges with initial vertex $v$ will be denoted $\operatorname{outdeg}(v)$, and the number of edges with final vertex $v$ will be denoted $\operatorname{indeg}(v)$. We say $D$ is Eulerian if it is connected, and at each vertex $v\in V(D)$, $\operatorname{indeg}(v)=\operatorname{outdeg}(v)$. A planar Eulerian digraph is called an alternating dimap if its planar embedding is such that, reading cyclically around each vertex, the edges alternate between incoming and outgoing.
@@ 1.2. Special alternating links from planar Eulerian digraphs
Special alternating links are associated to planar bipartite graphs $G$ or, equivalently, to $G$’s planar dual Eulerian digraph $D$. The edges of $D$ are oriented so that $D$ is an alternating dimap. See Figure 1.
We follow the construction for special alternating links presented by Juhász, Kálmán and Rasmussen [8] and by Kálmán and Murakami [9] to associate a positive special alternating link $L_{G}$ to a planar bipartite graph $G$. Let $M(G)$ be the medial graph of $G$: the vertex set of $M(G)$ is the set $\{v_{e}\ |\ e\in\operatorname{E}(G)\}$, and two vertices $v_{e}$ and $v_{e^{\prime}}$ of $M(G)$ are connected by an edge if the edges $e$ and $e^{\prime}$ are consecutive in the boundary of a face of $G$. We think of a particular planar drawing of $M(G)$ here: the midpoints of the edges of the planar drawing of $G$ are the vertices of $M(G)$. Thinking of $M(G)$ as a flattening of a link, there are two ways to choose under and overcrossings at each vertex of $M(G)$ to make it into an alternating link $L_{G}$. We select the over and undercrossings and orient $L_{G}$ so that each crossing is positive. This procedure yields a positive special alternating link.
More specifically, the orientation of $D$ will determine precisely how to orient $M(G)$ and select the over or undercrossings in order to generate $L_{G}$. We partition the regions of $M(G)$ corresponding to vertices of $G$ into color classes $A$ and $B$ such that whenever an edge of $D$ meets a vertex of $M(G)$, we pass a region in $A$ on the left and a region in $B$ on the right when traveling along the edge in the direction of its orientation. To obtain the associated positive special alternating link $L_{G}$, we orient the edges of $M(G)$ such that the boundary of every region in $A$ (resp. $B$) is a counterclockwise (resp. clockwise) cycle, then select over or undercrossings at each vertex of $M(G)$ such that all crossings of $L_{G}$ are positive. Any positive special alternating link arises from such a construction. Figure 1 shows an example.
@@ 1.3. The BEST Theorem for Eulerian digraphs
Recall that an Eulerian tour of $D$ is a directed cycle that uses each edge of $D$ exactly once. An oriented spanning tree rooted at $r\in V(D)$ is a spanning tree of $D$ such that for each vertex $v\in V(D)$, there is a unique directed path from $v$ to $r$.
@@
A $k$-spanning tree rooted at $r\in V(D)$ is a spanning tree of $D$ such that reversing the orientation of exactly $k$ of its edges yields an oriented spanning tree rooted at $r$. With this terminology, 00-spanning trees coincide with oriented spanning trees. Analogously, an arborescence rooted at $r$ is a spanning tree such that for every $v\in V(D)$, there is a unique directed path from $r$ to $v$, i.e. arborescences of $D$ are precisely $\{|V(D)|-1\}$-spanning trees.
@@
[16] For any $k\in\mathbb{N}$, the number of $k$-spanning trees of an Eulerian digraph $D$ is independent of the choice of root. Moreover, the sequence $c_{0}(D,r)$, normal-$\ldots$, $c_{|V(D)|-1}(D,r)$ is palindromic. Equivalently: $c_{k}(D,r)=c_{|V(D)|-1-k}(D,r)$ for all $k\in\{0,1,\ldots,|V(D)|-1\}$.
We will provide a new proof of Theorem 1.3 which will help us establish connections between the Alexander polynomials of special alternating links and volumes of root polytopes. This connection is outlined in Section 1.5.
@@ 1.4. The Alexander polynomial of an Eulerian digraph
Murasugi and Stoimenow [16] define the Alexander polynomial of an Eulerian digraph $D$ (and arbitrary $r\in V(D)$) to be:
@@
The name is justified by Murasugi and Stoimenow’s theorem that for an alternating dimap $D$, the polynomial $P_{D}(t)$ equals the Alexander polynomial of a special alternating link associated to $D$. Recall that an alternating dimap is a planar Eulerian digraph where at each vertex, the edges alternate between coming in and going out as we move cyclically around the vertex.
@@
[16, Theorem 2] For an alternating dimap $D$, the polynomial $P_{D}(t)$ equals the Alexander polynomial $\Delta_{L_{G}}(-t)$ for the special alternating link $L_{G}$ associated to the planar dual $G$ of $D$.
@@
Previous work of the authors [5, Theorem 1.2] proves that for alternating dimaps $D$, the Alexander polynomial $\Delta_{L_{G}}(-t)$ for the special alternating link $L_{G}$ associated to the planar dual $G$ of $D$ is log-concave. In light of Theorem 1.4, this result can be rephrased as follows for the polynomial  $P_{D}$:
@@
[5, Theorem 1.2] For any alternating dimap $D$, the coefficients of the polynomial $P_{D}(t)$ form a log-concave sequence with no internal zeros.
@@ 1.5. The polynomial $P_{D}$ in terms of root polytopes
The right-hand side of equation (2) was previously studied by Li and Postnikov [12] in the context of computing volumes of slices of zonotopes. Our proof techniques for Theorem 1.7 are different from Li and Postnikov’s [12] approach. We note the following interesting corollary for Alexander polynomials of special alternating links:
@@
(3) $\Delta_{L_{G}}(-t)=\sum_{\begin{subarray}{c}H\subset G\\ H\text{ connected}\end{subarray}}Vol(Q_{H})(q-1)^{|E(H)|-|V(G)|+1}.$
@@ 1.6. On the Kauffman and Crowell state models of the Alexander polynomial
The Alexander polynomial was initially defined by James Alexander II as a determinant [1]. Crowell, in 1959, gave the first combinatorial state model for Alexander polynomials of alternating links [3]. In John Conway’s paper [2], the Conway polynomial was introduced which satisfies a skein relation and is related to the Alexander polynomial by a substitution. In 1982, Kauffman defined the notion of states of a link universe, obtained from a link diagram. Originally, these states were used to compute a polynomial that could be specialized to the Conway polynomial. Later, in the most recent edition of [11], released in 2006, Kauffman described a means of using this combinatorial state model to directly compute the Alexander polynomial of any link.
@@
Roadmap of the paper. In Section 2, we study the enumeration of $k$-spanning trees of Eulerian digraphs which we use in Section 3 to prove Theorem 1.7. In Section 4, we show that for alternating dimaps $D$, $P_{D}(t)$ is the Alexander polynomial of a special alternating link. The proof we give is inspired by Murasugi and Stoimenow’s proof but is written in a graph theoretic language. In Section 5, we discuss possible generalizations of Fox’s long-standing conjecture that the sequence of absolute values of coefficients of the Alexander polynomial of an alternating link is unimodal. We conclude in Section 6 by constructing an explicit weight preserving bijection between the Crowell and Kauffman states of a special alternating link.
@@ 2. On $k$-spanning trees of Eulerian digraphs
([16, Proposition 1]) For any $k\in\mathbb{N}$, the number of $k$-spanning trees of an Eulerian digraph $D$ is independent of the choice of root. Moreover, the sequence $c_{0}(D,r)$, normal-$\ldots$, $c_{|\operatorname{V}(D)|-1}(D,r)$ is palindromic. Equivalently: $c_{k}(D,r)=c_{|\operatorname{V}(D)|-1-k}(D,r)$ for all $k\in\{0,1,\ldots,|\operatorname{V}(D)|-1\}$.
@@
([19, Theorem 10.2], [22, 23]) Let $D$ be an Eulerian digraph, and fix any edge $e\in\operatorname{E}(D)$ such that $\operatorname{init}(e)=r$. Let $\epsilon(D,e)$ denote the number of Eulerian tours of $D$ beginning with $e$. Then
@@
([19, Corollary 10.3], [22, 23]) For any Eulerian digraph $D$, $c_{0}(D,r)$ is independent of the root $r\in\operatorname{V}(D)$.
@@
The second statement follows by observing that 00-spanning trees of $D$ (oriented spanning trees) are precisely $(|V(D)|-1)$-spanning trees (arborescences) of $D^{T}$ and vice-versa. ∎
We note that Corollary 2.2 also follows from [21, Theorem 4.9].
For an arbitrary digraph $D$ and subset $E^{\prime}\subset E(D)$, let $D\setminus E^{\prime}$ denote the graph obtained from $D$ by deleting the edges in $E^{\prime}$, and let $D/E^{\prime}$ denote the graph obtained from $D$ by contracting the edges in $E^{\prime}$. A minor of $D$ is any graph which can be obtained from $D$ by deleting and/or contracting edges.
@@
Let $D$ be an Eulerian digraph, and fix any $r\in V(D)$. Then
$c_{k}(D,r)=\sum_{i=0}^{k}(-1)^{i}\sum_{\begin{subarray}{c}\text{\emph{acyclic}}E^{\prime}\subset E(D)\\ |E^{\prime}|=k-i\end{subarray}}\binom{|V(D)|-1-(k-i)}{i}c_{0}(D/E^{\prime},r)$
@@
Observe that if $T$ is a 00-spanning tree in $D/E^{\prime}$ for some acyclic $E^{\prime}\subset E(D)$, then $T_{D}:=(\operatorname{V}(D),E(T)\cup E^{\prime})$ is an $i$-spanning tree in $D$ rooted at $r$ for some $0\leq i\leq k$. Similarly, if $T_{D}$ is an $i$-spanning tree in $D$ and $E_{\text{away}}$ is the set of edges of $T_{D}$ directed away from the root, then $T_{D}/E^{\prime}$ is a 00-spanning tree in $D/E^{\prime}$ for any acyclic $E^{\prime}\supset E_{\text{away}}$. We can thus write
$c_{k}(D,r)=\sum_{\begin{subarray}{c}\text{acyclic }E^{\prime}\subset E(D)\\ |E^{\prime}|=k\end{subarray}}c_{0}(D/E^{\prime},r)-\sum_{\begin{subarray}{c}\text{acyclic }E^{\prime}\subset E(D)\\ |E^{\prime}|=k\end{subarray}}c_{<k}(D,r)$
where $c_{<k}(D,r)$ is the number of 00-spanning trees $T$ in $D/E^{\prime}$ such that there exists $e\in E^{\prime}$ where $E(T)\cup\{e\}$ forms a 0-spanning tree in $D/(E^{\prime}-\{e\})$. Using inclusion-exclusion, we rewrite this as follows:
$c_{k}(D,r)=\sum_{\begin{subarray}{c}\text{acyclic }E^{\prime}\subset E(D)\\ |E^{\prime}|=k\end{subarray}}c_{0}(D/E^{\prime},r)-\sum_{\begin{subarray}{c}\text{acyclic }E^{\prime}\subset E(D)\\ |E^{\prime}|=k\end{subarray}}\sum_{i=1}^{k}\sum_{\begin{subarray}{c}F\subset E^{\prime}\\ |F|=i\end{subarray}}(-1)^{i-1}c_{0}(D/(E^{\prime}-F),r).$
Observe that for any subset $F^{\prime}\subset E(D)$ with $|F^{\prime}|=k-i$ and any 00-spanning tree $T$ in $D/F^{\prime}$, contracting any $i$ edges of $T$ produces a 00-spanning tree of $D/E^{\prime}$ for some $E^{\prime}$ with $|E^{\prime}|=k$. Relabeling $F^{\prime}:=E^{\prime}-F$ in the above equation thus yields
$c_{k}(D,r)=\sum_{\begin{subarray}{c}\text{acyclic }E^{\prime}\subset E(D)\\ |E^{\prime}|=k\end{subarray}}c_{0}(D/E^{\prime},r)-\sum_{i=1}^{k}(-1)^{i-1}\sum_{\begin{subarray}{c}\text{acyclic }F^{\prime}\subset E(D)\\ |F^{\prime}|=k-i\end{subarray}}\binom{|V(D/F^{\prime})|-1}{i}c_{0}(D/F^{\prime},r).$
Since $|V(D/F^{\prime})|=|V(D)|-(k-i)$, we can relabel and combine the sums to obtain
$c_{k}(D,r)=\sum_{i=0}^{k}(-1)^{i}\sum_{\begin{subarray}{c}\text{acyclic }E^{\prime}\subset E(D)\\ |E^{\prime}|=k-i\end{subarray}}\binom{|V(D)|-1-(k-i)}{i}c_{0}(D/E^{\prime},r).$
@@
For any Eulerian digraph $D$, $c_{k}(D,r)$ is independent of the root $r$, and the sequence $c_{0}(D,r)$, normal-$\ldots$, $c_{|V(D)|-1}(D,r)$ is palindromic.
@@
Since $k$-spanning trees of $D$ are precisely $(|V(D)|-1-k)$-spanning trees of $D^{T}$, this follows immediately from Theorems 1.2 and 2.4. ∎

### v3

@@ Abstract.
The Alexander polynomial (1928) is the first polynomial invariant of links devised to help distinguish links up to isotopy. Fox’s conjecture (1962) – stating that the absolute values of the coefficients of the Alexander polynomial for any alternating link are trapezoidal – was settled for special alternating links by the present authors (2023); Kálmán, the second author, and Postnikov gave an alternative proof (2025). The present paper is a study of the special combinatorial and discrete geometric properties that Alexander polynomials of special alternating links possess along with a generalization to all Eulerian graphs, introduced by Murasugi and Stoimenow (2003). We prove that the Murasugi and Stoimenow generalized Alexander polynomials can be expressed in terms of volumes of root polytopes of unimodular matrices. The latter generalizes a result regarding the Alexander polynomials of special alternating links that follows by putting together the work of Li and Postnikov (2013) and Kálmán, the second author, and Postnikov (2025). Furthermore, we conjecture a generalization of Fox’s conjecture to the generalized Alexander polynomials of Murasugi and Stoimenow and bijectively relate two longstanding combinatorial models for the Alexander polynomials of special alternating links: Crowell’s state model (1959) and Kauffman’s state model (1982, 2006).
Karola Mészáros received support from NSF Grants DMS-1847284 and  DMS-2348676. Elena S. Hafner received support from NSF Grants DMS-1847284 and DMS-2402150. Alexander Vidinas received support from NSF Grant DMS-1847284.
@@ 1. Introduction
A tantalizing conjecture of Fox from 1962 [5] regarding the coefficients of the Alexander polynomials of alternating links remains stubbornly open to this day. It states that the absolute values of the coefficients of the Alexander polynomial of any alternating link are trapezoidal. Recall that a sequence $a_{1},\ldots,a_{n}$ is trapezoidal if
$a_{1}<a_{2}<\cdots<a_{k}=a_{k+1}=\cdots=a_{m}>a_{m+1}>\cdots>a_{n}$
for some $1\leq k\leq m\leq n$. For any alternating link $L$, [4, 15, 16] show that the Alexander polynomial $\Delta_{L}(t)$ can be normalized such that $\Delta_{L}(-t)\in\mathbb{Z}_{\geq 0}[t]$. In particular, the coefficients of $\Delta_{L}(t)$ have alternating signs, so Fox’s conjecture can be restated in the following way:
@@
[5] Let $L$ be an alternating link. Then the coefficients of $\Delta_{L}(-t)$ form a trapezoidal sequence.
Stoimenow [22] strengthened Fox’s conjecture to log-concavity without internal zeros. Fox’s conjecture remains open in general although some special cases have been settled by Hartley [7] for two-bridge knots, Murasugi [17] for a family of alternating algebraic links, and Ozsváth and Szabó [19] for the case of genus $2$ alternating knots. Jong [8] also confirmed that Fox’s conjecture holds for genus $2$ alternating knots. The authors of the current paper resolved Fox’s conjecture for special alternating links [6] in 2023; Azarpendar, Juhász, and Kálmán [2] built on these results to resolve Fox’s conjecture for certain diagrammatic Murasugi sums of special alternating links in 2024; while Kálmán, the second author, and Postnikov [10] gave an alternative proof of the special alternating case with a matroidal view in 2025.
@@ 1.1. Notation and graph theory definitions
For any graph $G$, let $\operatorname{V}(G)$ denote the vertex set of $G$ and $\operatorname{E}(G)$ the edge set of $G$. For a digraph $D$, we will denote the initial vertex of any $e\in\operatorname{E}(D)$ by $\operatorname{init}(e)$ and the final vertex of $e$ by $\operatorname{fin}(e)$. The number of edges with initial vertex $v$ will be denoted $\operatorname{outdeg}(v)$, and the number of edges with final vertex $v$ will be denoted $\operatorname{indeg}(v)$. We say $D$ is Eulerian if it is connected, and at each vertex $v\in\operatorname{V}(D)$, $\operatorname{indeg}(v)=\operatorname{outdeg}(v)$. A planar Eulerian digraph is called an alternating dimap if its planar embedding is such that, reading cyclically around each vertex, the edges alternate between incoming and outgoing.
@@ 1.2. Special alternating links from planar Eulerian digraphs
There are two types of crossings in a link diagram: positive and negative.
or
Crossings of the form on the left are called positive crossings and those of the form on the right are called negative crossings.
A special alternating link is any link which arises from the following construction, as presented by Juhász, Kálmán and Rasmussen [9] and by Kálmán and Murakami [11]. Let $G$ be a planar bipartite graph, and let $D$ be its planar dual. Observing that $D$ is Eulerian, we orient its edges so that $D$ is an alternating dimap. Let $M(G)$ be the medial graph of $G$: the vertex set of $M(G)$ is the set $\{v_{e}\ |\ e\in\operatorname{E}(G)\}$, and two vertices $v_{e}$ and $v_{e^{\prime}}$ of $M(G)$ are connected by an edge if the edges $e$ and $e^{\prime}$ are consecutive in the boundary of a face of $G$. We think of a particular planar drawing of $M(G)$ here: the midpoints of the edges of the planar drawing of $G$ are the vertices of $M(G)$. Thinking of $M(G)$ as a flattening of a link, the orientation of $D$ will determine how we orient $M(G)$. Partition the regions of $M(G)$ corresponding to vertices of $G$ into color classes $A$ and $B$ such that whenever an edge of $D$ meets a vertex of $M(G)$, we pass a region in $A$ on the left and a region in $B$ on the right when traveling along the edge in the direction of its orientation. We orient the edges of $M(G)$ such that the boundary of every region in $A$ (resp. $B$) is a counterclockwise (resp. clockwise) cycle and replace the vertices of $M(G)$ with over- and under-crossings such that either every crossing is positive or every crossing is negative. The case where every crossing is positive is called a positive special alternating link, and the case where every crossing is negative is a negative special alternating link. We denote by $L_{G}$ the positive special alternating link arising from a planar bipartite graph $G$ in this way. See Figure 1 for an example.
@@ 1.3. The BEST Theorem for Eulerian digraphs
Recall that an Eulerian tour of $D$ is a directed cycle that uses each edge of $D$ exactly once. An oriented spanning tree rooted at $r\in\operatorname{V}(D)$ is a spanning tree of $D$ such that for each vertex $v\in\operatorname{V}(D)$, there is a unique directed path from $v$ to $r$.
@@
A $k$-spanning tree rooted at $r\in\operatorname{V}(D)$ is a spanning tree of $D$ such that reversing the orientation of exactly $k$ of its edges yields an oriented spanning tree rooted at $r$. With this terminology, 00-spanning trees coincide with oriented spanning trees. Analogously, an arborescence rooted at $r$ is a spanning tree such that for every $v\in\operatorname{V}(D)$, there is a unique directed path from $r$ to $v$, i.e. arborescences of $D$ are precisely $\{|\operatorname{V}(D)|-1\}$-spanning trees.
@@
[18] For any $k\in\mathbb{N}$, the number of $k$-spanning trees of an Eulerian digraph $D$ is independent of the choice of root. Moreover, the sequence $c_{0}(D,r)$, $\ldots$, $c_{|\operatorname{V}(D)|-1}(D,r)$ is palindromic. Equivalently: $c_{k}(D,r)=c_{|\operatorname{V}(D)|-1-k}(D,r)$ for all $k\in\{0,1,\ldots,|\operatorname{V}(D)|-1\}$.
In Section 1.5, we will provide a new proof of Theorem 1.3 which will help us connect the Alexander polynomials of special alternating links to volumes of root polytopes.
@@ 1.4. The Alexander polynomial of an Eulerian digraph
Murasugi and Stoimenow [18] define the Alexander polynomial of an Eulerian digraph $D$ (and arbitrary $r\in\operatorname{V}(D)$) to be:
@@
The name is justified by Murasugi and Stoimenow’s theorem that for an alternating dimap $D$, the polynomial $P_{D}(t)$ equals the Alexander polynomial of a special alternating link associated to $D$. Recall that an alternating dimap is a planar Eulerian digraph such that, reading cyclically around each vertex, the edges alternate between coming in and going out.
@@
[18, Theorem 2] For an alternating dimap $D$, the polynomial $P_{D}(t)$ equals the Alexander polynomial $\Delta_{L_{G}}(-t)$ for the special alternating link $L_{G}$ associated to the planar dual $G$ of $D$, up to multiplication by ±$\pm t^{k}$ for some integer  $k$.
@@
Previous work of the authors [6, Theorem 1.2] proves that for alternating dimaps $D$, the Alexander polynomial $\Delta_{L_{G}}(-t)$ for the special alternating link $L_{G}$ associated to the planar dual $G$ of $D$ is log-concave. In light of Theorem 1.4, this result can be rephrased as follows for the polynomial  $P_{D}$:
@@
[6, Theorem 1.2] For any alternating dimap $D$, the coefficients of the polynomial $P_{D}(t)$ form a log-concave sequence with no internal zeros.
@@ 1.5. The polynomial $P_{D}$ in terms of root polytopes
Here $Q_{H}$ is the root polytope of the matrix $H$ (see Definition 3.3). An immediate corollary of the above theorem is:
@@
(3) $\Delta_{L_{G}}(-t)\sim\sum_{\begin{subarray}{c}H\subset G\\ H\text{ connected}\end{subarray}}Vol(Q_{H})(t-1)^{|\operatorname{E}(H)|-|\operatorname{V}(G)|+1}.$
We note that Corollary 1.8 can also be deduced putting together the work of Li and Postnikov [14] and Kálmán, Mészáros, and Postnikov [10]. Indeed, the work of Li and Postnikov [14] was the original inspiration for our Theorem 1.7. The right-hand side of equation (3) was introduced by Li and Postnikov [14] in the context of computing volumes of slices of zonotopes. They proved the equation:
(4) $f_{M_{\vec{G}}}(t)=\sum_{\begin{subarray}{c}H\subset G\\ H\text{ connected}\end{subarray}}Vol(Q_{H})(t-1)^{|\operatorname{E}(H)|-|\operatorname{V}(G)|+1}$
for a polynomial $f_{M_{\vec{G}}}(t)$ associated to the oriented graphic matroid of a planar bipartite graph with orientation $\vec{G}$ where the edges are directed from one bipartition to the other. The details of the definition of this polynomial can be found in [14] or in [10]. In turn, Kálmán, Mészáros, and Postnikov [10] prove in their Theorem 5.5 that:
(5) $\Delta_{L_{G}}(-t)\sim f_{M_{\vec{G}}}(t).$
Putting equations (4) and (5) together, one readily obtains Corollary 1.8.
@@ 1.6. On the Kauffman and Crowell state models of the Alexander polynomial
The Alexander polynomial was initially defined by James Alexander II as a determinant [1]. Crowell, in 1959, gave the first combinatorial state model for Alexander polynomials of alternating links [4]. In John Conway’s paper [3], he introduces the Conway polynomial which satisfies a skein relation and is related to the Alexander polynomial by a substitution. In 1982, Kauffman defined the notion of states of a link universe, obtained from a link diagram. Originally, these states were used to compute a polynomial that could be specialized to the Conway polynomial. Later, in the most recent edition of [13], released in 2006, Kauffman described a way of using this combinatorial state model to directly compute the Alexander polynomial of any link.
@@
Roadmap of the paper. In Section 2, we study the enumeration of $k$-spanning trees of Eulerian digraphs which we use in Section 3 to prove Theorem 1.7. In Section 4, we show that for alternating dimaps $D$, $P_{D}(t)$ is the Alexander polynomial of a special alternating link. The proof we give is inspired by Murasugi and Stoimenow’s proof but is written in a graph theoretic language. In Section 5, we discuss possible generalizations of Fox’s long-standing conjecture that the sequence of absolute values of coefficients of the Alexander polynomial of an alternating link is trapezoidal. We conclude in Section 6 by constructing an explicit weight-preserving bijection between the Crowell and Kauffman states of any special alternating link.
@@ 2. On $k$-spanning trees of Eulerian digraphs
([18, Proposition 1]) For any $k\in\mathbb{N}$, the number of $k$-spanning trees of an Eulerian digraph $D$ is independent of the choice of root. Moreover, the sequence $c_{0}(D,r)$, $\ldots$, $c_{|\operatorname{V}(D)|-1}(D,r)$ is palindromic. Equivalently: $c_{k}(D,r)=c_{|\operatorname{V}(D)|-1-k}(D,r)$ for all $k\in\{0,1,\ldots,|\operatorname{V}(D)|-1\}$.
@@
([21, Theorem 10.2], [24, 25]) Let $D$ be an Eulerian digraph, and fix any edge $e\in\operatorname{E}(D)$ such that $\operatorname{init}(e)=r$. Let $\epsilon(D,e)$ denote the number of Eulerian tours of $D$ beginning with $e$. Then
@@
([21, Corollary 10.3], [24, 25]) For any Eulerian digraph $D$, $c_{0}(D,r)$ is independent of the root $r\in\operatorname{V}(D)$.
@@
The second statement follows by observing that 00-spanning trees of $D$ (oriented spanning trees) are precisely $(|\operatorname{V}(D)|-1)$-spanning trees (arborescences) of $D^{T}$ and vice-versa. ∎
We note that Corollary 2.2 also follows from [23, Theorem 4.9].
For an arbitrary digraph $D$ and subset $E^{\prime}\subset\operatorname{E}(D)$, let $D\setminus E^{\prime}$ denote the graph obtained from $D$ by deleting the edges in $E^{\prime}$, and let $D/E^{\prime}$ denote the graph obtained from $D$ by contracting the edges in $E^{\prime}$. A minor of $D$ is any graph which can be obtained from $D$ by deleting and/or contracting edges.
@@
Let $D$ be an Eulerian digraph, and fix any $r\in\operatorname{V}(D)$. Then
$c_{k}(D,r)=\sum_{i=0}^{k}(-1)^{i}\sum_{\begin{subarray}{c}\text{\emph{cycle-free} }E^{\prime}\subset\operatorname{E}(D)\\ |E^{\prime}|=k-i\end{subarray}}\binom{|\operatorname{V}(D)|-1-(k-i)}{i}c_{0}(D/E^{\prime},r)$
@@
Observe that if $T$ is a 00-spanning tree in $D/E^{\prime}$ for some cycle-free $E^{\prime}\subset\operatorname{E}(D)$ with $|E^{\prime}|=k$, then $T_{D}:=(\operatorname{V}(D),\operatorname{E}(T)\cup E^{\prime})$ is an $i$-spanning tree in $D$ rooted at $r$ for some $0\leq i\leq k$. Similarly, if $T_{D}$ is an $i$-spanning tree in $D$ and $\operatorname{E}_{\text{away}}(T_{D})$ is the set of edges of $T_{D}$ directed away from the root, then $T_{D}/E^{\prime}$ is a 00-spanning tree in $D/E^{\prime}$ for any cycle-free $E^{\prime}\supset\operatorname{E}_{\text{away}}(T_{D}/E^{\prime})$. We can thus write
$c_{k}(D,r)=\sum_{\begin{subarray}{c}\text{cycle-free }E^{\prime}\subset\operatorname{E}(D)\\ |E^{\prime}|=k\end{subarray}}c_{0}(D/E^{\prime},r)-\sum_{\begin{subarray}{c}\text{cycle-free }E^{\prime}\subset\operatorname{E}(D)\\ |E^{\prime}|=k\end{subarray}}c_{<k}(D,E^{\prime},r)$
where $c_{<k}(D,E^{\prime},r)$ is the number of 00-spanning trees $T$ in $D/E^{\prime}$ such that there exists $e\in E^{\prime}$ where $\operatorname{E}(T)\cup\{e\}$ forms a 0-spanning tree in $D/(E^{\prime}-\{e\})$. Using inclusion-exclusion, we rewrite this as follows:
$c_{k}(D,r)=\sum_{\begin{subarray}{c}\text{cycle-free }E^{\prime}\subset\operatorname{E}(D)\\ |E^{\prime}|=k\end{subarray}}c_{0}(D/E^{\prime},r)-\sum_{\begin{subarray}{c}\text{cycle-free }E^{\prime}\subset\operatorname{E}(D)\\ |E^{\prime}|=k\end{subarray}}\sum_{i=1}^{k}\sum_{\begin{subarray}{c}F\subset E^{\prime}\\ |F|=i\end{subarray}}(-1)^{i-1}c_{0}(D/(E^{\prime}-F),r).$
Observe that for any subset $F^{\prime}\subset\operatorname{E}(D)$ with $|F^{\prime}|=k-i$ and any 00-spanning tree $T$ in $D/F^{\prime}$, contracting any $i$ edges of $T$ produces a 00-spanning tree of $D/E^{\prime}$ for some $E^{\prime}$ with $|E^{\prime}|=k$. Relabeling $F^{\prime}:=E^{\prime}-F$ in the above equation thus yields
$c_{k}(D,r)=\sum_{\begin{subarray}{c}\text{cycle-free }E^{\prime}\subset\operatorname{E}(D)\\ |E^{\prime}|=k\end{subarray}}c_{0}(D/E^{\prime},r)-\sum_{i=1}^{k}(-1)^{i-1}\sum_{\begin{subarray}{c}\text{cycle-free }F^{\prime}\subset\operatorname{E}(D)\\ |F^{\prime}|=k-i\end{subarray}}\binom{|\operatorname{V}(D/F^{\prime})|-1}{i}c_{0}(D/F^{\prime},r).$
Since $|\operatorname{V}(D/F^{\prime})|=|\operatorname{V}(D)|-(k-i)$, we can relabel and combine the sums to obtain
$c_{k}(D,r)=\sum_{i=0}^{k}(-1)^{i}\sum_{\begin{subarray}{c}\text{cycle-free }E^{\prime}\subset\operatorname{E}(D)\\ |E^{\prime}|=k-i\end{subarray}}\binom{|\operatorname{V}(D)|-1-(k-i)}{i}c_{0}(D/E^{\prime},r).$
@@
([18, Proposition 1]) For any Eulerian digraph $D$, $c_{k}(D,r)$ is independent of the root $r$, and the sequence $c_{0}(D,r)$, $\ldots$, $c_{|\operatorname{V}(D)|-1}(D,r)$ is palindromic.
@@
Since $k$-spanning trees of $D$ are precisely $(|\operatorname{V}(D)|-1-k)$-spanning trees of $D^{T}$, this follows immediately from Theorems 1.2 and 2.4. ∎